# Exploring the Potentials of Spiking Neural Networks for Image Deraining

### v2

@@ Introduction
Spiking Neural Networks (SNNs) have recently emerged as highly efficient and biologically plausible alternatives to traditional Artificial Neural Networks (ANNs) and Deep Neural Networks (DNNs) with significant improvements in energy efficiency and computational simplicity. Recent studies have demonstrated the promising potential of SNNs across high-level computer vision tasks, including image classification (deng2022temporal), object detection (luo2024integer) and semantic segmentation (patel2021spiking), with effective sequential information processing and representation.
Given the remarkable efficiency and success in high-level visual tasks, a natural question arises: can SNNs also excel in low-level vision tasks? In this paper, we focus specifically on image deraining, a fundamental low-level vision task. Addressing this challenge requires answering a fundamental Question and overcoming a critical Obstacle. The Question is: what distinctive features do SNNs represent in the image deraining context? Unlike convolutional layers, which aggregate local information using weighted filters, the conventional spiking neuron, LIF (Leaky Integrate-and-Fire), integrates discrete spikes over time to trigger neuronal firing events (maass1997networks). Such temporal accumulation differs fundamentally from the spatially continuous convolutional operation, potentially leading to distinct feature extraction mechanisms. Although (song2024learning; su2025bridge; liang2025spikederain; xiao2025spiking) introduce SNNs to vision restoration, the existing works have not addressed this differentiation at the feature representation level, a gap particularly crucial for low-level vision tasks.
@@
To overcome the identified Obstacle, we draw inspiration from neurophysiological findings showing that neurons in the inferotemporal (IT) cortex of the ventral visual pathway integrate stimuli across spatially extended receptive fields to recognise complex object features (kobatake1994neuronal). Motivated by this mechanism, we leverage the cumulative firing behaviour of Leaky Integrate-and-Fire (LIF) neurons to provide visual awareness. However, due to the intrinsic firing-resetting mechanism of LIF neurons, such binary accumulation alone fails to represent rich visual features. To address this, we propose the Visual LIF (VLIF), combining the above insights with Integrate-and-Fire dynamics enhanced by continuous membrane potentials, thus equipping LIF neurons with spatially meaningful receptive fields without introducing additional parameters.
@@
Single image deraining has rapidly advanced with the application of deep learning, where initial convolutional neural network (CNN)-based methods typically approached deraining as pixel-wise regression tasks (jiang2020mspfn; zamir2021MPRNet). More recent methods have employed multi-stage or multi-scale CNN architectures and self-attention mechanisms to improve rain removal performance by capturing intricate patterns and modelling long-range dependencies (Zamir2022Restormer). Despite their effectiveness, these CNN and attention-based approaches often incur heavy computational costs, limiting deployment on resource-constrained devices. Very recently, a different paradigm was explored with spiking neural networks (song2024learning). The SNN-based approach opens a new direction for efficient rain removal on low-power devices.
@@
Spiking Neural Networks (SNNs) offer energy-efficient computation through sparse, event-driven processing, making them promising alternatives to conventional ANNs. SNNs are typically implemented via ANN-to-SNN conversion or direct training. The former maps a trained ANN to an SNN by replacing activations with spiking units (cao2015spiking_ann; diehl2015fast_ann_relu), but often requires long simulation time and suffers from performance degradation. In contrast, direct training with surrogate gradients enables end-to-end optimisation and achieves competitive accuracy with much fewer timesteps and lower latency (neftci2019surrogate). In computer vision, SNNs have been applied to high-level tasks such as image classification (hu2021ms-resnet), object detection (luo2024integer) and segmentation (patel2021spiking). However, their use in pixel-level low-level tasks (such as image deraining) remains rare. We explore this under-investigated area and demonstrate that SNNs can be effectively adapted for dense image regression.
@@
Patch-to-Time Reorganisation. Given the input feature $X\in\mathbb{R}^{T\times C\times H\times W}$, we first apply a pixel unshuffle (shi2016real) with downsampling factor $r$, downscale $X$ to $X^{\prime}\in\mathbb{R}^{T\times C\cdot r^{2}\times\frac{H}{r}\times\frac{W}{r}}$, which splits $X$ into $r^{2}$ non-overlapping spatial patches. Next, we reshape and permute the tensor by bringing the patch dimension into the temporal axis to have $X^{\prime\prime}\in\mathbb{R}^{T\cdot r^{2}\times C\times\frac{H}{r}\times\frac{W}{r}}$. This transformation effectively converts local spatial patches into temporally ordered sequences, which allows the neuron to integrate local visual context over the temporal axis. Therefore, low-intensity pixels within high-response regions are more likely to fire, thus enhancing spatial continuity in rain structure encoding:
@@
While VLIF fires based on spatial context, the spatial accumulation carries greater structural variability and finer-grained information. In such cases, binary spikes are insufficient to express continuous intensity differences. To address this, we consider I-LIF (luo2024integer), an integer-valued spiking neuron that generates discrete spike counts for better representation. However, similar to the observation in Spike2Former (lei2025spike2former), I-LIF leads to training instability. In our task, this manifests as integer overflow in the output images. We argue that the potential reason is that the large integer values accumulated by I-LIF lead to corrupted or saturated reconstruction results, which fail to represent finer-grained features.
To mitigate this, we adopt virtual timestep normalisation (lei2025spike2former), to bound the spike values and stabilise gradient flow, named NI-LIF, by scaling spike magnitudes with $D$ and apply clipping to bound $x$ to $[x,min,max]$:
@@
Following the spiking stage, we apply a temporal attention  (yao2023attention) to obtain the attended spike features with the shape ${r^{2}}T,C,H/r,W/r]$. We denote this output as $S_{\text{attn}}$, which is subsequently used for output compression.
@@
(ding2025rethinking) finds that the output of the vanilla LIF is poorly informative at the first timestep and only achieves decent performance after integrating subsequent temporal steps, as the LIF with tt=11 suffers from strong membrane potential suppression due to the zero-initialised hidden state. As a result, input features may fail to exceed the firing threshold at tt=11 (Fig. 2b), even with relatively high intensity. This behaviour limits LIF to represent fine-grained spatial details, which is critical in pixel-level restoration tasks such as image deraining. In contrast, VLIF normalises spike outputs and encourages early activation, yielding significantly denser responses. As shown in Fig. 3, LIF achieves only 16.46% activation, while VLIF reaches 58.94% activation ($3.6\times$ improvement) and maintains 25.4% fire rate even for the lowest input range [0.002, 0.114], where LIF is unable to fire. This showcases VLIF’s strength in preserving weak signals and reconstructing fine details through spatio-temporal reorganisation and multi-level quantisation.
@@
We then aggregate all components with a residual connection, followed by a convolution and batch normalisation (zheng2021tdbn) to refine the features:
@@
Additionally, a multi-dimensional attention module (MDA) (yao2023attention) enhances spatio-temporal and channel interactions. Then the final SDEM output is given:
@@
Inspired by the fact that multi-scale modelling at the shallow level can enhance representational capacity (cui2023image), we introduce a lightweight multi-scale module, SMU, which leverages the proposed SDEM and a VLIF-T, to incorporate fine-grained rain cues from both local and extended neighbourhoods with minimal computational overhead, offering an efficient multi-scale enhancement mechanism in highest-resolution stages of both encoder and decoder.
As illustrated in Fig. 4(c), SMU first receives input feature $\in\mathbb{R}^{T\times C\times H\times W}. xx$ then is processed by the proposed SDEM, followed by a VLIF-T neuron that further downsamples the spatial resolution to $[H/r,W/r]$ via patch-to-time reorganisation. The downsampled feature is then passed through two stacked Spiking Units (SU), each composed of a LIF spiking neuron, batch normalisation, convolution, and a multi-dimensional attention (MDA) module (yao2023attention). Two SUs further extract high-frequency representations at lower spatial resolutions. The features are subsequently upsampled back to $[T,C,H,W]$ using bilinear interpolation. To retain shallow spatial information, a residual connection bypasses intermediate layers, with merged features refined by an MDA block before the next step.
@@
To comprehensively evaluate the performance, we conduct experiments on five widely used synthetic rain datasets (Rain12 (li2016GMM), Rain200L (yang2017JORDER), Rain200H (yang2017JORDER), Rain1200 (zhang2018diddata)) and the Real-World rain dataset (ren2019PReNet). All datasets are used following the original train/test splits (shown in the supplementary material with training iterations). We evaluate model performance for synthetic data on two widely used metrics, Peak Signal to Noise Ratio (PSNR) (hore2010PSNR) and Structural Similarity (SSIM) (wang2004SSIM), shown in Tab. 1 (reproduced from (song2024learning)). We retrain ESDNet based on their official setting, as their pre-trained models are inaccessible. Perception-based Image Quality Evaluator (PIQE) (venkatanath2015PIQE) and Blind/Referenceless Image Spatial QUality Evaluator BRISQUE (mittal2012no) are used for the real-world data, shown in Tab. 2(a). Following (hu2021spikingresnet; song2024learning), energy is calculated with: 12.5pJ per FLOP, 77fJ per SOP (synaptic operation), and 3.7pJ per spike-based Sign operation (calculated using energy per spike). All ablation studies are conducted on RainH200 for 30K iterations.
@@
To further validate the task-oriented behaviour and generalisability of spiking neurons in image restoration, we train our model on the EUVP (islam2020fast) Underwater Image Enhancement (UIE) dataset, where the goal is to correct large-scale low-frequency degradation from light scattering and colour shifts. As shown in Tab. 2(d), our method surpasses recent UIE approaches (liu2022twin; ma2022wavelet; zhao2024toward) while using the least computation. Consistent with our frequency-domain analysis (Fig. 6(c)), LIF neurons act as low-frequency indicators in UIE, highlighting the task-adaptive nature of spiking neurons in image restoration.

### v3

@@ Introduction
Spiking Neural Networks (SNNs) have recently emerged as highly efficient and biologically plausible alternatives to traditional Artificial Neural Networks (ANNs) and Deep Neural Networks (DNNs) with significant improvements in energy efficiency and computational simplicity. Recent studies have demonstrated the promising potential of SNNs across high-level computer vision tasks, including image classification (Deng et al. 2022), object detection (Luo et al. 2024) and semantic segmentation (Patel et al. 2021), with effective sequential information processing and representation.
Given the remarkable efficiency and success in high-level visual tasks, a natural question arises: can SNNs also excel in low-level vision tasks? In this paper, we focus specifically on image deraining, a fundamental low-level vision task. Addressing this challenge requires answering a fundamental Question and overcoming a critical Obstacle. The Question is: what distinctive features do SNNs represent in the image deraining context? Unlike convolutional layers, which aggregate local information using weighted filters, the conventional spiking neuron, LIF (Leaky Integrate-and-Fire), integrates discrete spikes over time to trigger neuronal firing events (Maass 1997). Such temporal accumulation differs fundamentally from the spatially continuous convolutional operation, potentially leading to distinct feature extraction mechanisms. Although (Song et al. 2024; Su et al. 2025; Liang et al. 2025; Xiao et al. 2025) introduce SNNs to vision restoration, the existing works have not addressed this differentiation at the feature representation level, a gap particularly crucial for low-level vision tasks.
@@
To overcome the identified Obstacle, we draw inspiration from neurophysiological findings showing that neurons in the inferotemporal (IT) cortex of the ventral visual pathway integrate stimuli across spatially extended receptive fields to recognise complex object features (Kobatake and Tanaka 1994). Motivated by this mechanism, we leverage the cumulative firing behaviour of Leaky Integrate-and-Fire (LIF) neurons to provide visual awareness. However, due to the intrinsic firing-resetting mechanism of LIF neurons, such binary accumulation alone fails to represent rich visual features. To address this, we propose the Visual LIF (VLIF), combining the above insights with Integrate-and-Fire dynamics enhanced by continuous membrane potentials, thus equipping LIF neurons with spatially meaningful receptive fields without introducing additional parameters.
@@
Single image deraining has rapidly advanced with the application of deep learning, where initial convolutional neural network (CNN)-based methods typically approached deraining as pixel-wise regression tasks (Jiang et al. 2020; Zamir et al. 2021). More recent methods have employed multi-stage or multi-scale CNN architectures and self-attention mechanisms to improve rain removal performance by capturing intricate patterns and modelling long-range dependencies (Zamir et al. 2022). Despite their effectiveness, these CNN and attention-based approaches often incur heavy computational costs, limiting deployment on resource-constrained devices. Very recently, a different paradigm was explored with spiking neural networks (Song et al. 2024). The SNN-based approach opens a new direction for efficient rain removal on low-power devices.
@@
Spiking Neural Networks (SNNs) offer energy-efficient computation through sparse, event-driven processing, making them promising alternatives to conventional ANNs. SNNs are typically implemented via ANN-to-SNN conversion or direct training. The former maps a trained ANN to an SNN by replacing activations with spiking units (Cao et al. 2015; Diehl et al. 2015), but often requires long simulation time and suffers from performance degradation. In contrast, direct training with surrogate gradients enables end-to-end optimisation and achieves competitive accuracy with much fewer timesteps and lower latency (Neftci et al. 2019). In computer vision, SNNs have been applied to high-level tasks such as image classification (Hu et al. 2021b), object detection (Luo et al. 2024) and segmentation (Patel et al. 2021). However, their use in pixel-level low-level tasks (such as image deraining) remains rare. We explore this under-investigated area and demonstrate that SNNs can be effectively adapted for dense image regression.
@@
Patch-to-Time Reorganisation. Given the input feature $X\in\mathbb{R}^{T\times C\times H\times W}$, we first apply a pixel unshuffle (Shi et al. 2016) with downsampling factor $r$, downscale $X$ to $X^{\prime}\in\mathbb{R}^{T\times C\cdot r^{2}\times\frac{H}{r}\times\frac{W}{r}}$, which splits $X$ into $r^{2}$ non-overlapping spatial patches. Next, we reshape and permute the tensor by bringing the patch dimension into the temporal axis to have $X^{\prime\prime}\in\mathbb{R}^{T\cdot r^{2}\times C\times\frac{H}{r}\times\frac{W}{r}}$. This transformation effectively converts local spatial patches into temporally ordered sequences, which allows the neuron to integrate local visual context over the temporal axis. Therefore, low-intensity pixels within high-response regions are more likely to fire, thus enhancing spatial continuity in rain structure encoding:
@@
While VLIF fires based on spatial context, the spatial accumulation carries greater structural variability and finer-grained information. In such cases, binary spikes are insufficient to express continuous intensity differences. To address this, we consider I-LIF (Luo et al. 2024), an integer-valued spiking neuron that generates discrete spike counts for better representation. However, similar to the observation in Spike2Former (Lei et al. 2025), I-LIF leads to training instability. In our task, this manifests as integer overflow in the output images. We argue that the potential reason is that the large integer values accumulated by I-LIF lead to corrupted or saturated reconstruction results, which fail to represent finer-grained features.
To mitigate this, we adopt virtual timestep normalisation (Lei et al. 2025), to bound the spike values and stabilise gradient flow, named NI-LIF, by scaling spike magnitudes with $D$ and apply clipping to bound $x$ to $[x,min,max]$:
@@
Following the spiking stage, we apply a temporal attention  (Yao et al. 2023) to obtain the attended spike features with the shape ${r^{2}}T,C,H/r,W/r]$. We denote this output as $S_{\text{attn}}$, which is subsequently used for output compression.
@@
(Ding et al. 2025) finds that the output of the vanilla LIF is poorly informative at the first timestep and only achieves decent performance after integrating subsequent temporal steps, as the LIF with tt=11 suffers from strong membrane potential suppression due to the zero-initialised hidden state. As a result, input features may fail to exceed the firing threshold at tt=11 (Fig. 2b), even with relatively high intensity. This behaviour limits LIF to represent fine-grained spatial details, which is critical in pixel-level restoration tasks such as image deraining. In contrast, VLIF normalises spike outputs and encourages early activation, yielding significantly denser responses. As shown in Fig. 3, LIF achieves only 16.46% activation, while VLIF reaches 58.94% activation ($3.6\times$ improvement) and maintains 25.4% fire rate even for the lowest input range [0.002, 0.114], where LIF is unable to fire. This showcases VLIF’s strength in preserving weak signals and reconstructing fine details through spatio-temporal reorganisation and multi-level quantisation.
@@
We then aggregate all components with a residual connection, followed by a convolution and batch normalisation (Zheng et al. 2021) to refine the features:
@@
Additionally, a multi-dimensional attention module (MDA) (Yao et al. 2023) enhances spatio-temporal and channel interactions. Then the final SDEM output is given:
@@
Inspired by the fact that multi-scale modelling at the shallow level can enhance representational capacity (Cui et al. 2023), we introduce a lightweight multi-scale module, SMU, which leverages the proposed SDEM and a VLIF-T, to incorporate fine-grained rain cues from both local and extended neighbourhoods with minimal computational overhead, offering an efficient multi-scale enhancement mechanism in highest-resolution stages of both encoder and decoder.
As illustrated in Fig. 4(c), SMU first receives input feature $\in\mathbb{R}^{T\times C\times H\times W}. xx$ then is processed by the proposed SDEM, followed by a VLIF-T neuron that further downsamples the spatial resolution to $[H/r,W/r]$ via patch-to-time reorganisation. The downsampled feature is then passed through two stacked Spiking Units (SU), each composed of a LIF spiking neuron, batch normalisation, convolution, and a multi-dimensional attention (MDA) module (Yao et al. 2023). Two SUs further extract high-frequency representations at lower spatial resolutions. The features are subsequently upsampled back to $[T,C,H,W]$ using bilinear interpolation. To retain shallow spatial information, a residual connection bypasses intermediate layers, with merged features refined by an MDA block before the next step.
@@
To comprehensively evaluate the performance, we conduct experiments on five widely used synthetic rain datasets (Rain12 (Li et al. 2016), Rain200L (Yang et al. 2017), Rain200H (Yang et al. 2017), Rain1200 (Zhang and Patel 2018)) and the Real-World rain dataset (Ren et al. 2019). All datasets are used following the original train/test splits (shown in the supplementary material with training iterations). We evaluate model performance for synthetic data on two widely used metrics, Peak Signal to Noise Ratio (PSNR) (Hore and Ziou 2010) and Structural Similarity (SSIM) (Wang et al. 2004), shown in Tab. 1 (reproduced from (Song et al. 2024)). We retrain ESDNet based on their official setting, as their pre-trained models are inaccessible. Perception-based Image Quality Evaluator (PIQE) (Venkatanath et al. 2015) and Blind/Referenceless Image Spatial QUality Evaluator BRISQUE (Mittal et al. 2012) are used for the real-world data, shown in Tab. 2(a). Following (Hu et al. 2021a; Song et al. 2024), energy is calculated with: 12.5pJ per FLOP, 77fJ per SOP (synaptic operation), and 3.7pJ per spike-based Sign operation (calculated using energy per spike). All ablation studies are conducted on RainH200 for 30K iterations.
@@
To further validate the task-oriented behaviour and generalisability of spiking neurons in image restoration, we train our model on the EUVP (Islam et al. 2020) Underwater Image Enhancement (UIE) dataset, where the goal is to correct large-scale low-frequency degradation from light scattering and colour shifts. As shown in Tab. 2(d), our method surpasses recent UIE approaches (Liu et al. 2022; Ma and Oh 2022; Zhao et al. 2024) while using the least computation. Consistent with our frequency-domain analysis (Fig. 6(c)), LIF neurons act as low-frequency indicators in UIE, highlighting the task-adaptive nature of spiking neurons in image restoration.